OpenStudy (anonymous):

What is the solution to 3 times the square root of 4 minus 2 times the square root of 4?

OpenStudy (anonymous):

\( \large Since~you~seem~new~here:\\\bf\Huge{~~\color{red}{\boxed{W}}\color{#FF9200}{\boxed{E}}\color{#FFDB00}{\boxed{L}}\color{#B6ff00}{\boxed{C}}\color{#00ff49}{\boxed{O}}\color{#00DBff}{\boxed{M}}\color{#B600ff}{\boxed{E}}\\~~~~~~~~~~~~~~~~~~\boxed{T}\boxed{O}\\~~~~~~~~~~~\color{#0092ff}{\boxed{O}}\color{#0092ff}{\boxed{P}}\color{#0092ff}{\boxed{E}}\color{#0092ff}{\boxed{N}}\color{#7cc517}{\boxed{S}}\color{#7cc517}{\boxed{T}}\color{#7cc517}{\boxed{U}}\color{#7cc517}{\boxed{D}}\color{#7cc517}{\boxed{Y}}\color{#7cc517}{\boxed{!}}} \\ \large\bf You~can~read~the~\underline{\href{ /code-of-conduct }{Code~of~Conduct}}~here. \\ \it \large Now~let's~try~to~find~an~answer~to~your~problem~together.\)

OpenStudy (anonymous):

That would equal to 2(: 3 * √4 = 6 2 * √4 = 4 6 - 4 = 2(:

OpenStudy (anonymous):

\[\Large 3\sqrt{4} - 2\sqrt{4}=\]

OpenStudy (anonymous):

did i read that right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@bwright12 do understand the answer of yuba?